# Mykola Khrypchenko

### v3

@@ Introduction
Let $n$ be a positive integer and $\mathcal{A}$ a variety of algebras with one $n$-linear operation [,,][$\cdot$,$\dots$,$\cdot$][ ⋅ , … , ⋅ ]. Consider two $n$-linear operations $[\cdot,\dots,\cdot]_{1}$ and $[\cdot,\dots,\cdot]_{2}$ on the same vector space $V$, such that the algebras $(V,[\cdot,\dots,\cdot]_{1})$ and $(V,[\cdot,\dots,\cdot]_{2})$ belong to $\mathcal{A}$. Then $[\cdot,\dots,\cdot]_{1}$ and $[\cdot,\dots,\cdot]_{2}$ are said to be compatible if $(V,[\cdot,\dots,\cdot]_{1}+[\cdot,\dots,\cdot]_{2})$ belongs to $\mathcal{A}$. In this case, the triple $(V,[\cdot,\dots,\cdot]_{1},[\cdot,\dots,\cdot]_{2})$ is called a compatible $\mathcal{A}$-algebra. For example, there are compatible associative algebras [6, 17], compatible Lie algebras [12], compatible pre-Lie algebras [1], compatible Leibniz algebras [16] and so on. If the signature of a variety $\mathcal{A}$ has several operations (of possibly different types), one can consider the compatibility with respect to some fixed operation, for example, compatible Poisson brackets [3], compatible Hom-Lie algebras [8], compatible Hom-Lie triple systems [21] and so on, or with respect to all the operations at the same time as in [20, Definition A].
@@
In Section 1 we specify the definitions of $\sigma$-matching ($\sigma\in S_{2}$), interchangeable and total compatibility to the case of a pair of associative products and give examples showing the difference between the compatibilities. As often happens, nilpotent algebras are a good source of such examples. The case of unital products is considered in Section 2 (see Propositions 2.2 and 2.6) as a motivation for Section 3, where we prove some general results on idempotent algebras (see Propositions 3.1 and 3.2) and proceed to more specific classes of such algebras: the semigroup algebras of rectangular bands and algebras with enough idempotents (the latter class includes some algebras of infinite matrices, path algebras and infinite direct sums of unital algebras). The descriptions of $\sigma$-matching and interchangeable products on these algebras are given in Propositions 3.3, 3.8, 3.18 and 3.23. Section 4 is devoted to free non-unital associative algebras. We first prove a general result valid for all algebras without zero divisors (Proposition 4.1) and then treat separately the cases of non-commutative and commutative non-unital free algebras (see Propositions 4.2, 4.6, 4.9 and 4.12).
@@ 1.1 Compatible structures
Let us consider the following particular cases of 1.
@@
    $(a\cdot_{1}b)\cdot_{2}c=a\cdot_{\sigma(1)}(b\cdot_{\sigma(2)}c)\text{ and }(a\cdot_{2}b)\cdot_{1}c=a\cdot_{\sigma(2)}(b\cdot_{\sigma(1)}c),$ (2)
@@
$(a\cdot_{1}b)\cdot_{2}c=(a\cdot_{2}b)\cdot_{1}c\text{ and }a\cdot_{1}(b\cdot_{2}c)=a\cdot_{2}(b\cdot_{1}c).$ (4)
@@
    $\cdot_{1}$ and $\cdot_{2}$ are interchangeable and at least one of the two equalities 2 holds for some $\sigma\in S_{2}$;
@@
    $\cdot_{1}$ and $\cdot_{2}$ are $\sigma$-matching for some $\sigma\in S_{2}$ and at least one of the two equalities 4 holds;
@@
    $\cdot_{1}$ and $\cdot_{2}$ are $\sigma_{1}$-matching for some $\sigma_{1}\in S_{2}$ and at least one of the two equalities 2 holds for $\sigma\neq\sigma_{1}$;
If $\mathrm{char}(K)\neq 2$, then each of the conditions (i), (ii), (iii) and (iv) is also equivalent to
@@
The product * on $A$ given by $e_{1}*e_{1}=e_{1}$ and $e_{1}*e_{2}=e_{2}$ is clearly associative (observe that $(A,*)$ is isomorphic to the direct sum of a $1$-dimensional algebra with zero multiplication and the subalgebra of $M_{2}(K)$ generated by the matrix units $\{E_{11},E_{12}\}$). Then $(e_{i}*e_{j})\cdot e_{k}\neq 0\Leftrightarrow e_{i}\cdot(e_{j}*e_{k})\neq 0\Leftrightarrow(i,j,k)=(1,1,2)$, in which case $(e_{i}*e_{j})\cdot e_{k}=e_{i}\cdot(e_{j}*e_{k})=e_{3}$. Moreover, $(A\cdot A)*A=A*(A\cdot A)=\{0\}$. Thus, the product * is $(12)$-matching with $\cdot$, but it is neither $\mathrm{id}$-matching nor interchangeable with $\cdot$ because $(A*A)\cdot A\neq\{0\}$.
@@
Given two structures $*_{1}$ and $*_{2}$ on $A$ that are compatible with $\cdot$, we say that $*_{1}$ and $*_{2}$ are isomorphic, if there exists an automorphism $\phi$ of $(A,\cdot)$ such that
@@ 2 $\sigma$-matching and interchangeable structures on unital associative algebras
Although unital associative algebras are a subclass of algebras with enough idempotents considered in Section 3, we begin with this particular case to motivate the choice of the context for Section 3. We also introduce here some notions that will be used below.
@@
is interchangeable with $\cdot$. Moreover, if $\cdot$ is associative, then $*_{\varphi}$ is also associative and totally compatible with $\cdot$.
@@
The product of the form 5 is said to be determined by $\varphi\in\Gamma(A)$.
@@
(i)$\Rightarrow$(iv). Let * be $(12)$-matching with $\cdot$. Then for all $a\in A$ we have
$11*a=(1*a)\cdot 1=1\cdot(a*1)=a*$ 1 .
@@
(ii)$\Rightarrow$(iv). Let * be interchangeable with $\cdot$. Then for all $a\in A$ we have
$1*a=(1\cdot 1)*a=(1*1)\cdot a,\text{ and }a*1=a*(1\cdot 1)=a\cdot(1*1).$
Now, $11*a=(1*a)\cdot 1=(1\cdot a)*1=a*$ 1, whence $(1*1)\cdot a=a\cdot(1*1)$, i.e. $1*1\in C(A)$. Thus,
@@
Let $x$ be a non-central element of $A$. Then $\cdot_{x}$ is $\mathrm{id}$-matching with $\cdot$ by Lemma 2.1. If $\cdot_{x}$ was totally compatible with $\cdot$, then by Proposition 2.6 there would be $c\in C(A)$ such that $\cdot_{x}=\cdot_{c}$, whence $x=1\cdot_{x}1=1\cdot_{c}1=c$, a contradiction. ∎
@@
Whenever $\Gamma(A)=K$, there are only two non-isomorphic structures of the form 5: the original product $\cdot$ and the zero product.
For, in this case $a*_{\varphi}b=\lambda(a\cdot b)$ for some $\lambda\in K$, so, if $\lambda\neq 0$, the map $\lambda{}^{-1}\mathrm{id}$ is an isomorphism between $\cdot$ and $*_{\varphi}$.
@@
For non-unital algebras the result of Proposition 2.6 may be false. For example, if $A$ is an algebra with zero multiplication, then any associative product * on $A$ is totally compatible with $\cdot$ since all the mixed monomials in 3 are zero. However, all the mutations of $\cdot$ and all the products on $A$ determined by elements of $\Gamma(A)$ are zero.
@@ 3 $\sigma$-matching and interchangeable structures on certain idempotent associative algebras
In this section we are going to see to which extent the results of Section 2 generalize to several classes of associative algebras that are in some sense close to being unital. Recall that a (not-necessarily associative) algebra $(A,\cdot)$ is said to be idempotent [^1] In the context of Lie algebras one prefers to use the term “perfect” rather than “idempotent”. if $A\cdot A=A$.
@@
We only need to prove (i)$\Rightarrow$(ii). Let * be interchangeable with $\cdot$. Then, for all $a,b,c\in A$ with $b=b_{1}\cdot b_{2}$ we have
@@
Since $A\cdot A=A$, then by linearity $a\cdot(b*c)=(a*b)\cdot c$ for all $a,b,c\in A$. Thus, * is totally compatible with $\cdot$ by Remark 1.3(ii). ∎
It is natural to ask if one can replace interchangeable products by $\sigma$-matching ones in Proposition 3.1. For $\sigma=\mathrm{id}$ the answer is “no” even in the case of unital algebras, as we saw in Corollary 2.7. For $\sigma=(12)$ there are classes of idempotent algebras containing the unital ones for which the answer is positive and those for which it is negative. We first point out a class admitting the positive answer.
@@
We only need to prove (i)$\Rightarrow$(ii). Assume that $A$ has a left unit $e$. Let * be $(12)$-matching with $\cdot$. Then for all $a,b,c\in A$ we have
@@
It follows that * is totally compatible with $\cdot$ by Remark 1.3(iv). If $A$ has a right unit $e$, then we symmetrically have
@@ 3.1 The semigroup algebra of a rectangular band
which equals $\mu_{ij,kl}e_{il}$ for some $\mu_{ij,kl}\in K$ by 6. Moreover, for any $(p,q)\in I\times J$ we have
@@
whence $\mu_{ij,kl}=\mu_{pj,kl}$. Thus, $\mu_{ij,kl}$ does not depend on $i$ and $l$, so denoting it by $\lambda_{jk}$, we get 7.
Conversely, let * be given by 7. Then for all $(i,j),(k,l),(p,q)\in I\times J$ we have
@@
The product 7 is a mutation of $\cdot$ if and only if $\lambda_{jk}$ does not depend on $j$ and $k$.
Indeed, given $a=\sum_{p,q}a_{pq}e_{pq}$ with $a_{pq}\in K$, we have $e_{ij}\cdot_{a}e_{kl}=e_{ij}\cdot a\cdot e_{kl}=\lambda e_{il}$, where $\lambda=\sum_{p,q}a_{pq}$. Conversely, if $\lambda_{jk}=\lambda$ for all $j$ and $k$ in 7, then $e_{ij}*e_{kl}=\lambda e_{il}=e_{ij}\cdot_{a}\cdot e_{kl}$, where $a=\lambda e_{pq}$ for some fixed arbitrary $(p,q)\in I\times J$.
@@
Assuming that $a\cdot e_{ij}=e_{ij}\cdot a$, for all $(p,q)\in I\times J$ by 9 we have
@@
Since $e_{ij}\cdot(e_{ij}*e_{ij})=(e_{ij}*e_{ij})\cdot e_{ij}$, by Lemma 3.6 there are $\lambda_{ij}\in K$ and $r_{ij}\in\operatorname{Ann}(A)$, such that $e_{ij}*e_{ij}=\lambda_{ij}e_{ij}+r_{ij}$. Now, on the one hand $(e_{ij}*e_{ij})\cdot e_{il}=(\lambda_{ij}e_{ij}+r_{ij})\cdot e_{il}=\lambda_{ij}e_{il}$, and on the other hand using 8 we have
@@
imply $\lambda_{ij}=\lambda_{kj}$ for all $i,k\in I$ and $j\in J$. Thus, $\lambda_{ij}=\lambda_{kl}$ for all $(i,j),(k,l)\in I\times J$, and 10 follows. ∎
@@
By Lemmas 3.5 and 3.7 any $(12)$-matching structure on $(A,\cdot)$ has the form 11. Conversely, let * be a product on $A$ given by 11. Then for all $(i,j),(k,l),(p,q)\in I\times J$ we have $(e_{ij}*e_{kl})\cdot e_{pq}=(\lambda e_{il}+r_{il})\cdot e_{pq}=\lambda e_{iq}$ and $e_{ij}\cdot(e_{kl}*e_{pq})=e_{ij}\cdot(\lambda e_{kq}+r_{kq})=\lambda e_{iq}$. Moreover, $(e_{ij}\cdot e_{kl})*e_{pq}=e_{il}*e_{pq}=\lambda e_{iq}+r_{iq}$ and $e_{ij}*(e_{kl}\cdot e_{pq})=e_{ij}*e_{kq}=\lambda e_{iq}+r_{iq}$. ∎
@@
Let * be totally compatible with $\cdot$. In particular, * is $(12)$-matching with $\cdot$. Then for all $(i,j)\in I\times J$ by 11 we have $(e_{ij}*e_{ij})\cdot e_{ij}=\lambda e_{ij}$, while $e_{ij}*(e_{ij}\cdot e_{ij})=e_{ij}*e_{ij}=\lambda e_{ij}+r_{ij}$. Thus, $r_{ij}=0$ for all $(i,j)\in I\times J$ in 11, and we get 12. Conversely, the product 12 is $*_{\varphi}$, where $\varphi=\lambda\mathrm{id}\in\Gamma(A)$, so * is totally compatible with $\cdot$ by Lemma 2.3. ∎
@@
Let $\varphi\in\Gamma(A)$ and consider the totally compatible product $*_{\varphi}$ on $A$ as in Lemma 2.3. By Corollary 3.9 there exists $\lambda\in K$ such that $\varphi(e_{ij})=\varphi(e_{ij}\cdot e_{ij})=e_{ij}*_{\varphi}e_{ij}=\lambda e_{ij}$ for all $(i,j)\in I\times J$. By linearity, $\varphi=\lambda\mathrm{id}$. ∎
@@
This is a consequence of Propositions 3.1, 2.3 and 3.9. ∎
In general, the annihilator part of 11 may be non-trivial, which makes it difficult to characterize explicitly the associativity of 11 in terms of $\lambda$ and $r_{il}$. So, we just give an example of such an associative product below, which also provides a $(12)$-matching structure on $A$ that is not totally compatible with $\cdot$, showing that one cannot replace interchangeable structures by $(12)$-matching ones in Proposition 3.1.
@@
Then $(A*A)*A=A*(A*A)=\{0\}$, so that $(A,*)$ is associative. Observe that * is of the form 11 for $\lambda=0$ and $r_{il}=e_{11}-e_{12}-e_{21}+e_{22}\in\operatorname{Ann}(A)$ for $(i,l)=(1,2)$ and $r_{il}=0$ otherwise. Thus, * is $(12)$-matching with $\cdot$, but it is not totally compatible with $\cdot$ by Corollary 3.9.
@@
In view of Propositions 3.1 and 3.11 we only need to show (i)$\Leftrightarrow$(iii). If $I=\{i_{0}\}$, then $e_{i_{0}j}\cdot e_{i_{0}k}=e_{i_{0}k}$ for all $j,k\in J$, whence $e_{i_{0}j}\cdot a=a$ for all $a\in A$, so that each $e_{i_{0}j}$ is a left unit of $A$, and the result follows by Proposition 3.2 (or from Corollary 3.9, since the existence of a left unit implies $\operatorname{Ann}(A)=\{0\}$). Whenever $J=\{j_{0}\}$, each $e_{ij_{0}}$ is a right unit of $A$, and we apply Proposition 3.2 again. ∎

### v4

@@ Introduction
Let $n$ be a positive integer and $\mathcal{A}$ a variety of algebras with one $n$-linear operation $\cdot$$\dots,\cdot$]. Consider two $n$-linear operations $\cdot,\dots,\cdot]_{1}$ and $\cdot,\dots,\cdot]_{2}$ on the same vector space $V$, such that the algebras $(V,[\cdot,\dots,\cdot]_{1})$ and $(V,[\cdot,\dots,\cdot]_{2})$ belong to $\mathcal{A}$. Then $\cdot,\dots,\cdot]_{1}$ and $\cdot,\dots,\cdot]_{2}$ are said to be compatible if $(V,[\cdot,\dots,\cdot]_{1}+[\cdot,\dots,\cdot]_{2})$ belongs to $\mathcal{A}$. In this case, the triple $(V,[\cdot,\dots,\cdot]_{1},[\cdot,\dots,\cdot]_{2})$ is called a compatible $\mathcal{A}$-algebra. For example, there are compatible associative algebras [6, 17], compatible Lie algebras [12], compatible pre-Lie algebras [1], compatible Leibniz algebras [16] and so on. If the signature of a variety $\mathcal{A}$ has several operations (of possibly different types), one can consider the compatibility with respect to some fixed operation, for example, compatible Poisson brackets [3], compatible Hom-Lie algebras [8], compatible Hom-Lie triple systems [21] and so on, or with respect to all the operations at the same time as in [20, Definition A].
@@
In Section˜1 we specify the definitions of $\sigma$-matching ($\sigma\in S_{2}$), interchangeable and total compatibility to the case of a pair of associative products and give examples showing the difference between the compatibilities. As often happens, nilpotent algebras are a good source of such examples. The case of unital products is considered in Section˜2 (see Propositions˜2.2 and 2.6) as a motivation for Section˜3, where we prove some general results on idempotent algebras (see Propositions˜3.1 and 3.2) and proceed to more specific classes of such algebras: the semigroup algebras of rectangular bands and algebras with enough idempotents (the latter class includes some algebras of infinite matrices, path algebras and infinite direct sums of unital algebras). The descriptions of $\sigma$-matching and interchangeable products on these algebras are given in Propositions˜3.3, 3.8, 3.18 and 3.23. Section˜4 is devoted to free non-unital associative algebras. We first prove a general result valid for all algebras without zero divisors (Proposition˜4.1) and then treat separately the cases of non-commutative and commutative non-unital free algebras (see Propositions˜4.2, 4.6, 4.9 and 4.12).
@@ 1.1 Compatible structures
Let us consider the following particular cases of ˜1.
@@
[…]
@@
[…]
@@
    $\cdot_{1}$ and $\cdot_{2}$ are interchangeable and at least one of the two equalities ˜2 holds for some $\sigma\in S_{2}$;
@@
    $\cdot_{1}$ and $\cdot_{2}$ are $\sigma$-matching for some $\sigma\in S_{2}$ and at least one of the two equalities ˜4 holds;
@@
    $\cdot_{1}$ and $\cdot_{2}$ are $\sigma_{1}$-matching for some $\sigma_{1}\in S_{2}$ and at least one of the two equalities ˜2 holds for $\sigma\neq\sigma_{1}$;
If $\mathrm{char}(K)\neq 2$, then each of the conditions ˜(i), (ii), (iii) and (iv) is also equivalent to
@@
The product * on $A$ given by $e_{1}*e_{1}=e_{1}$ and $e_{1}*e_{2}=e_{2}$ is clearly associative (observe that (A,)(A,*) is isomorphic to the direct sum of a $1$-dimensional algebra with zero multiplication and the subalgebra of $M_{2}(K)$ generated by the matrix units $\{E_{11},E_{12}\}$). Then $(e_{i}*e_{j})\cdot e_{k}\neq 0\Leftrightarrow e_{i}\cdot(e_{j}*e_{k})\neq 0\Leftrightarrow(i,j,k)=(1,1,2)$, in which case $(e_{i}*e_{j})\cdot e_{k}=e_{i}\cdot(e_{j}*e_{k})=e_{3}$. Moreover, $(A\cdot A)*A=A*(A\cdot A)=\{0\}$. Thus, the product * is $(12)$-matching with $\cdot$, but it is neither $\mathrm{id}$-matching nor interchangeable with $\cdot$ because $(A*A)\cdot A\neq\{0\}$.
@@
Given two structures $_{1}$ and $_{2}$ on $A$ that are compatible with $\cdot$, we say that $_{1}$ and $_{2}$ are isomorphic, if there exists an automorphism $\phi$ of $(A,\cdot)$ such that
@@ 2 $\sigma$-matching and interchangeable structures on unital associative algebras
Although unital associative algebras are a subclass of algebras with enough idempotents considered in Section˜3, we begin with this particular case to motivate the choice of the context for Section˜3. We also introduce here some notions that will be used below.
@@
is interchangeable with $\cdot$. Moreover, if $\cdot$ is associative, then $_{\varphi}$ is also associative and totally compatible with $\cdot$.
@@
The product of the form ˜5 is said to be determined by $\varphi\in\Gamma(A)$.
@@
[…]
$1*a=(1*a)\cdot 1=1\cdot(a*1)=a*1.$
@@
[…]
Now, $1*a=(1*a)\cdot 1=(1\cdot a)*1=a*1$, whence $(1*1)\cdot a=a\cdot(1*1)$, i.e. $1*1\in C(A)$. Thus,
@@
Let $x$ be a non-central element of $A$. Then $\cdot_{x}$ is $\mathrm{id}$-matching with $\cdot$ by Lemma˜2.1. If $\cdot_{x}$ was totally compatible with $\cdot$, then by Proposition˜2.6 there would be $c\in C(A)$ such that $\cdot_{x}=\cdot_{c}$, whence $x=1\cdot_{x}1=1\cdot_{c}1=c$, a contradiction. ∎
@@
Whenever $\Gamma(A)=K$, there are only two non-isomorphic structures of the form ˜5: the original product $\cdot$ and the zero product.
For, in this case $a*_{\varphi}b=\lambda(a\cdot b)$ for some $\lambda\in K$, so, if $\lambda\neq 0$, the map $\lambda{}^{-1}\mathrm{id}$ is an isomorphism between $\cdot$ and $_{\varphi}$.
@@
For non-unital algebras the result of Proposition˜2.6 may be false. For example, if $A$ is an algebra with zero multiplication, then any associative product * on $A$ is totally compatible with $\cdot$ since all the mixed monomials in ˜3 are zero. However, all the mutations of $\cdot$ and all the products on $A$ determined by elements of $\Gamma(A)$ are zero.
@@ 3 $\sigma$-matching and interchangeable structures on certain idempotent associative algebras
In this section we are going to see to which extent the results of Section˜2 generalize to several classes of associative algebras that are in some sense close to being unital. Recall that a (not-necessarily associative) algebra $(A,\cdot)$ is said to be idempotent [^1] In the context of Lie algebras one prefers to use the term “perfect” rather than “idempotent”. if $A\cdot A=A$.
@@
[…]
@@
Since $A\cdot A=A$, then by linearity $a\cdot(b*c)=(a*b)\cdot c$ for all $a,b,c\in A$. Thus, * is totally compatible with $\cdot$ by Remark˜1.3˜(ii). ∎
It is natural to ask if one can replace interchangeable products by $\sigma$-matching ones in Proposition˜3.1. For $\sigma=\mathrm{id}$ the answer is “no” even in the case of unital algebras, as we saw in Corollary˜2.7. For $\sigma=(12)$ there are classes of idempotent algebras containing the unital ones for which the answer is positive and those for which it is negative. We first point out a class admitting the positive answer.
@@
[…]
@@
It follows that * is totally compatible with $\cdot$ by Remark˜1.3˜(iv). If $A$ has a right unit $e$, then we symmetrically have
@@ 3.1 The semigroup algebra of a rectangular band
which equals $\mu_{ij,kl}e_{il}$ for some $\mu_{ij,kl}\in K$ by ˜6. Moreover, for any $(p,q)\in I\times J$ we have
@@
whence $\mu_{ij,kl}=\mu_{pj,kl}$. Thus, $\mu_{ij,kl}$ does not depend on $i$ and $l$, so denoting it by $\lambda_{jk}$, we get ˜7.
Conversely, let * be given by ˜7. Then for all $(i,j),(k,l),(p,q)\in I\times J$ we have
@@
The product ˜7 is a mutation of $\cdot$ if and only if $\lambda_{jk}$ does not depend on $j$ and $k$.
Indeed, given $a=\sum_{p,q}a_{pq}e_{pq}$ with $a_{pq}\in K$, we have $e_{ij}\cdot_{a}e_{kl}=e_{ij}\cdot a\cdot e_{kl}=\lambda e_{il}$, where $\lambda=\sum_{p,q}a_{pq}$. Conversely, if $\lambda_{jk}=\lambda$ for all $j$ and $k$ in ˜7, then $e_{ij}*e_{kl}=\lambda e_{il}=e_{ij}\cdot_{a}\cdot e_{kl}$, where $a=\lambda e_{pq}$ for some fixed arbitrary $(p,q)\in I\times J$.
@@
Assuming that $a\cdot e_{ij}=e_{ij}\cdot a$, for all $(p,q)\in I\times J$ by ˜9 we have
@@
Since $e_{ij}\cdot(e_{ij}*e_{ij})=(e_{ij}*e_{ij})\cdot e_{ij}$, by Lemma˜3.6 there are $\lambda_{ij}\in K$ and $r_{ij}\in\operatorname{Ann}(A)$, such that $e_{ij}*e_{ij}=\lambda_{ij}e_{ij}+r_{ij}$. Now, on the one hand $(e_{ij}*e_{ij})\cdot e_{il}=(\lambda_{ij}e_{ij}+r_{ij})\cdot e_{il}=\lambda_{ij}e_{il}$, and on the other hand using ˜8 we have
@@
imply $\lambda_{ij}=\lambda_{kj}$ for all $i,k\in I$ and $j\in J$. Thus, $\lambda_{ij}=\lambda_{kl}$ for all $(i,j),(k,l)\in I\times J$, and ˜10 follows. ∎
@@
By Lemmas˜3.5 and 3.7 any $(12)$-matching structure on $(A,\cdot)$ has the form ˜11. Conversely, let * be a product on $A$ given by ˜11. Then for all $(i,j),(k,l),(p,q)\in I\times J$ we have $(e_{ij}*e_{kl})\cdot e_{pq}=(\lambda e_{il}+r_{il})\cdot e_{pq}=\lambda e_{iq}$ and $e_{ij}\cdot(e_{kl}*e_{pq})=e_{ij}\cdot(\lambda e_{kq}+r_{kq})=\lambda e_{iq}$. Moreover, $(e_{ij}\cdot e_{kl})*e_{pq}=e_{il}*e_{pq}=\lambda e_{iq}+r_{iq}$ and $e_{ij}*(e_{kl}\cdot e_{pq})=e_{ij}*e_{kq}=\lambda e_{iq}+r_{iq}$. ∎
@@
Let * be totally compatible with $\cdot$. In particular, * is $(12)$-matching with $\cdot$. Then for all $(i,j)\in I\times J$ by ˜11 we have $(e_{ij}*e_{ij})\cdot e_{ij}=\lambda e_{ij}$, while $e_{ij}*(e_{ij}\cdot e_{ij})=e_{ij}*e_{ij}=\lambda e_{ij}+r_{ij}$. Thus, $r_{ij}=0$ for all $(i,j)\in I\times J$ in ˜11, and we get ˜12. Conversely, the product ˜12 is $_{\varphi}$, where $\varphi=\lambda\mathrm{id}\in\Gamma(A)$, so * is totally compatible with $\cdot$ by Lemma˜2.3. ∎
@@
Let $\varphi\in\Gamma(A)$ and consider the totally compatible product $_{\varphi}$ on $A$ as in Lemma˜2.3. By Corollary˜3.9 there exists $\lambda\in K$ such that $\varphi(e_{ij})=\varphi(e_{ij}\cdot e_{ij})=e_{ij}*_{\varphi}e_{ij}=\lambda e_{ij}$ for all $(i,j)\in I\times J$. By linearity, $\varphi=\lambda\mathrm{id}$. ∎
@@
This is a consequence of Propositions˜3.1, 2.3 and 3.9. ∎
In general, the annihilator part of ˜11 may be non-trivial, which makes it difficult to characterize explicitly the associativity of ˜11 in terms of $\lambda$ and $r_{il}$. So, we just give an example of such an associative product below, which also provides a $(12)$-matching structure on $A$ that is not totally compatible with $\cdot$, showing that one cannot replace interchangeable structures by $(12)$-matching ones in Proposition˜3.1.
@@
Then $(A*A)*A=A*(A*A)=\{0\}$, so that (A,)(A,*) is associative. Observe that * is of the form ˜11 for $\lambda=0$ and $r_{il}=e_{11}-e_{12}-e_{21}+e_{22}\in\operatorname{Ann}(A)$ for $(i,l)=(1,2)$ and $r_{il}=0$ otherwise. Thus, * is $(12)$-matching with $\cdot$, but it is not totally compatible with $\cdot$ by Corollary˜3.9.
@@
[…]